# IDDM: Bridging Synthetic-to-Real Domain Gap from Physics-Guided Diffusion for Real-world Image Dehazing

### v1

@@ 3.1 Forward Process
$\small q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=N(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1}+\sqrt{\bar{\alpha}_{t}}\Delta\mathbf{h}_{t},\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{t-1}),$ (13)
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=N(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}(\mathbf{x}_{0}+\mathbf{h}_{t}),\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}_{t}}).$ (14)
In the end, haze of varying densities is combined with noise and fed into the model over $T$ steps. During the forward training process, $\mathbf{h}_{t}$ serves as a conditional vector and a haze component. This enables the model to learn how clear images and haze gradually combine, allowing it to remove haze during inference physically. Unlike existing methods, our forward process learns the distribution of $\mathbf{x}_{0}+\mathbf{h}_{t}$ rather than just $\mathbf{x}_{0}$. Therefore, learning the distribution of $\mathbf{h}_{t}$ becomes necessary to derive the distribution of $\mathbf{x}_{0}$. While $\mathbf{x}_{0}+\mathbf{h}_{t}$ shifts the data range slightly beyond $[-1,1]$, simple normalization corrects this deviation.
Figure 3: The flowchart of our method illustrates three stages of training. In stage 1, a denoising Unet is trained with the known $\mathbf{h}_{t}$, and then in stage 2, it is used to train the HtNet. In stage 3, IDDM is trained with the $\mathbf{\tilde{h}}_{t}$ estimated by HtNet.
By learning $\mathbf{h}_{t}$ at each time step $t$, which represents haze concentration at different depths, our model is able to progressively remove haze from a hazy image for $T$ times. This is particularly effective for distant haze, which is typically denser. Multiple stages of image dehazing are beneficial for removing distant haze, a challenge that previous methods often struggled to address.
3.2 Sampling Process
The determinisic implicit sampling process of IDDM can be derived by referring DDIM [41], shown in Fig. 2.
$\mathbf{x_{t-1}}=$ $\sqrt{\bar{\alpha}_{t-1}}\left(\mathbf{x_{0}+h_{t}}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\boldsymbol{\epsilon}_{\theta}^{(t)}+\sigma_{t}\boldsymbol{\epsilon}_{\theta}^{(t)}$ (15)
@@ 3.2 Sampling Process
where $\boldsymbol{\epsilon}_{\theta}^{(t)}=\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},\mathbf{h}_{t},\mathbf{\hat{x}},t\right)$. Then based on Eq. (15), computational cost can be reduced by utilizing only a subset of timestep indices, which is the same as DDIM [41]. Instead of considering all timesteps from $1$ to $T$, a subsequence $(\tau_{1},\tau_{2},\ldots,\tau_{s})$ is selected. This subsequence is determined by evenly spacing out indices from the complete sequence $(1,2,\ldots,T)$ using the formula $T_{i}=(i-1)\cdot\frac{T}{S}+1$, where $i$ ranges from $1$ to $\tau_{s}$ and $s$ is a parameter controlling the sampling rate. Therefore, for step $\tau_{1}<\tau_{2}$, the sampling process is given by
@@ 4 Method
4.1 Training Procedure
The training procedure is illustrated as shown in Fig. 3. In the forward process, we continuously introduce the effects of haze into the denoising Unet, thereby enhancing the noise estimator’s ability to adapt to haze conditions. Conversely, we also leverage haze to help the denoising Unet more accurately estimate noise in images taken in hazy environments. In the sampling process, we perform dehazing and denoising on $x_{t}$ simultaneously.
The training process can be divided into 3 stages. In stages 1 and 2, we do not utilize paired data (hazy and clear images) from known datasets because $\mathbf{h}_{t}$ cannot be directly obtained from such pairs. To address this issue, we generate both hazy images and $\mathbf{h}_{t}$ during the diffusion process, using their corresponding depth maps and clear images. We set the airlight $A\in[0.7,1.0]$, and the scattering coefficient $\beta\in[0.4,1.5]$. As $t$ progresses from 00 to $T$, a clear image gradually transforms into a hazy image. By randomly generating hazy images during training, we ensure both the randomness of the training process and the diversity of the data. In stage 3, we train the final model using paired data to ensure fair comparison with other methods.
At stage 1, we randomly generate the ground truth $\mathbf{h}_{t}$ at time step $t$ and hazy image $\hat{\mathbf{x}}$. A denoising Unet is trained with noise image $\mathbf{x}_{t}$ computed with Eq. (12), hazy image $\hat{\mathbf{x}}$, and $\mathbf{h}_{t}$, as shown in Fig. 3 (a),(c) and (d). Although the denoising Unet can compute the noise of images at distribution $q\left(\mathbf{x}_{t},\mathbf{x}_{0}\right)$ with the help of $\mathbf{h}_{t}$, the value of $\mathbf{h}_{t}$ remains unknown for a degraded image $\tilde{{\mathbf{x}}}$. Given that $\mathbf{h}_{t}$ is related to time $t$ and $\hat{\mathbf{x}}$, it is estimated with the assistance of the noise predictor. At stage 2, shown in Fig. 3 (e), we train a $\mathbf{h}_{t}$ estimator, denoted as ’HtNet’, with the model trained in stage 1. Here, the network structure of HtNet corresponds to the denoising Unet trained at stage 1 but with 1/3 input channels. The optimization at this stage has two objectives. Firstly, the predicted $\tilde{\mathbf{h}_{t}}$ should be close to the ground truth ${\mathbf{h}_{t}}$. Secondly, when the predicted $\tilde{\mathbf{h}_{t}}$ is fed into the denoising Unet, the estimated noise should closely match the noise predicted using the ground truth $\mathbf{h}_{t}$. Finally, at stage 3, the diffusion model is retrained with the HtNet. All experimental settings remain the same as in stage 1, except that $\mathbf{h}_{t}$ is now provided by HtNet.
Figure 4: Visual comparisons on real-world images from RTTS dataset [21] and Fattal’s dataset [10]. All models are trained on OTS dataset. Please zoom in for more details.
Training Dataset ITS OTS
@@ 4.1 Training Procedure
Dehamer [13] 29.3234 4.5819 45.7394 4.5998 33.8582 4.9134 46.5483 4.6256
DEA-Net [4] 29.7903 4.8192 45.813 4.6635 31.1197 4.9206 48.1424 4.6499
MB-TaylorFormer [35] 32.8450 4.9090 47.7943 4.6302 32.6807 4.8664 46.8968 4.6184
C2PNet [56] 33.7423 5.0546 48.5107 4.6739 34.2951 5.0336 48.6310 4.6399
PSD [3] N/A N/A N/A N/A 25.2394 3.8742 30.5665 4.3459
D4 [50] 33.206 5.8548 57.7313 3.7239 N/A N/A N/A N/A
OKNet [7] 33.1591 4.9664 49.1645 4.6558 35.1734 5.0858 49.1408 4.6548
IDDM 28.4986 4.2872 40.8716 4.8545 20.3332 3.4604 28.7864 4.8776
Table 1: The quantitative results of our method and the state-of-the-arts on real-world haze dataset RTTS. The best results are shown in red, and the second best results are shown in blue.
4.2 Loss Function
We optimize our models with a general way, and the loss function are concisely based on the training procedure. At stage 1, for a hazy image $\hat{\mathbf{x}}$, the noise predictor $\boldsymbol{\tilde{\epsilon}}_{\theta}$ is optimized by
$\mathcal{L}_{s_{1}}=\mathbb{E}_{t\sim[1,T],\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right)}\left\|\boldsymbol{\epsilon}_{t}-\boldsymbol{\tilde{\epsilon}_{\theta}}\left(\mathbf{x}_{t},\mathbf{h}_{t},\hat{\mathbf{x}},t\right)\right\|^{2}.$ (18)
Next, at stage 2, we adopt the denoising Unet $\boldsymbol{\tilde{\epsilon}_{{\theta}}}$ to optimize the $\mathbf{h}_{t}$ estimator, HtNet. Firstly, the output $\tilde{\mathbf{h}}_{t}$ should closely match its ground truth $\mathbf{h}_{t}$. Secondly, the noise predicted by $\boldsymbol{\tilde{\epsilon}_{{\theta}}}$ based on $\tilde{\mathbf{h}}_{t}$ should be close to that using $\mathbf{h}_{t}$. Therefore, we design a loss function to achieve the optimization objectives outlined above:
At stage 3, we also use MSE loss with the $\tilde{\mathbf{h}}_{t}$ estimated by HtNet to regularize the final denoisin Unet $\boldsymbol{\epsilon}_{{\theta}}$ by
$\mathcal{L}_{s_{3}}=\mathbb{E}_{t\sim[1,T],\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right)}\left\|\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{{\theta}}\left(\mathbf{x}_{t},\tilde{\mathbf{h}_{t}},\hat{\mathbf{x}},t\right)\right\|^{2}.$ (20)
@@ 5.1 Experimental Settings
Datasets: During training, we use ITS (13,990 indoor synthetic hazy image pairs) and OTS (72,135 outdoor synthetic hazy image pairs) datasets from RESIDE [21]. ITS depth maps come from NYU Depth Dataset V2 [33], while OTS provides its own depth maps. We evaluate our method against state-of-the-art techniques on RTTS [21] (4,322 real-world hazy images) and BEDDE [54] (209 paired real-world images) datasets, with additional visual comparisons using Fattal’s dataset [10].
Figure 5: Visual comparisons on real-world images from RTTS dataset [21]. All models are trained on ITS dataset. Please zoom in for more details.
Implementation Details: During the training stages, experiments are conducted on 4 NVIDIA GeForce RTX 4090 GPUs. We optimize our networks with Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.99$). The learning rate is set to 0.00001 and the batch size is set to 16. For data argumentation, we randomly flip images both horizontally and vertically and crop them to a size of 256*256. At stages 1 and 3, the diffusion denoising Unet is trained for 500,000 iterations, with the parameter $T$ is set as 1000. At stage 2, the $\mathbf{h}_{t}$ estimator is trained for 6w iterations.
Evaluation Metrics: For RTTS dataset [21], we employ four well-known non-reference metrics, NIQE [27], Brisque [28], PIQE [29] and NIMA [43] to evaluate perceptual quality of the real-world images. For the BEDDE dataset [54], we employ the evaluation metrics provided by the dataset’s authors: VI, RI, and VSI.
@@ 5.2 Results and Comparisons
Quantitative Comparison: To verify the robustness of domain generalization, we train the models on both the indoor synthetic hazy dataset [21] and the outdoor synthetic hazy dataset [21], and test them on two real-world datasets, RTTS [21] and BEDDE [54]. We compare our method with 7 state-of-the-arts approaches: namely Dehamer [13], DEA-Net [4], MB-TaylorFormer [35], C2PNet [56], PSD [3], D4 [50], and OKNET [7]. Among all of them, PSD [3] is trained solely on outdoor hazy images, and D4 [50] is solely trained on indoor hazy images. As shown in Table 1 and Table 2, IDDM achieves almost the best results on seven metrics across two datasets. Our IDDM achieves the best quantitative performance compared to the other approaches.
@@
(d) w/o Stage 3
(e) ours
Figure 6: Visual ablation results of the proposed IDDM.
Qualitative Comparison: As shown in Fig. 5 and Fig. 4, we compare the visual results of our model with previous state-of-the-art methods on the RTTS dataset [21] and Fattal’s data [10]. In Fig. 5, all models are trained on indoor synthetic dataset, while in Fig. 4, all models are trained on the outdoor synthetic dataset. As shown in Fig. 5, even when the training and test sets span two domains (indoor to outdoor, synthetic to real), IDDM still shows bright and clear dehazing performance. In contrast, the other methods tend to produce dark tones in the dehazed images, likely due to being trained exclusively on indoor hazy images. As shown in Fig. 4, we also include the classic fattal’s dataset [10] in addition to the RTTS dataset.
Trained on ITS Trained on OTS
[…]
Dehamer [13] 0.805 0.942 0.909 0.795 0.950 0.909
DEA-Net [4] 0.840 0.934 0.899 0.882 0.973 0.949
MB-TaylorFormer [35] 0.793 0.933 0.898 0.793 0.933 0.898
C2PNet [56] 0.795 0.913 0.874 0.880 0.970 0.948
D4/PSD [50, 3] 0.814 0.951 0.923 0.860 0.966 0.925
OKNet [7] 0.846 0.935 0.908 0.875 0.973 0.947
IDDM (Ours) 0.879 0.959 0.936 0.903 0.973 0.956
Table 2: Quantitative comparison on BEDDE dataset. Red and blue colors indicate the best and second-best results respectively.
@@
$\beta\in[0.1,1.0]$ 0.641 0.948 0.838 $T=300$ 0.669 0.905 0.865
$\beta\in[0.4,1.5]$ 0.879 0.959 0.936 $T=500$ 0.750 0.922 0.881
$\beta\in[1.0,2.0]$ 0.616 0.911 0.822 $T=1000$ 0.879 0.959 0.936
Table 4: Ablation studies on hyperparameters $\beta$ (scattering coefficient range) and $T$ (diffusion timesteps) evaluated on the BEDDE dataset. Models are trained on OTS.
@@
Table 5: Ablation studies on hyperparameter $A$ (atmospheric light range) evaluated on the BEDDE dataset. Models are trained on ITS.

### v3

@@ 3.3 The IDDM Forward Process
$\small q(\mathbf{x}_{t}|\mathbf{x}_{t-1},\mathbf{h}_{t})=N(\mathbf{x}_{t};\sqrt{\alpha_{t}}\mathbf{x}_{t-1}+\sqrt{\bar{\alpha}_{t}}\Delta\mathbf{h}_{t},\sqrt{1-\alpha_{t}}\boldsymbol{\mathbf{I}}),$ (9)
$q(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{h}_{t})=N(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}(\mathbf{x}_{0}+\mathbf{h}_{t}),\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\mathbf{I}}).$ (10)
In the end, haze of varying densities is combined with noise and fed into the model over $T$ steps. During the forward training process, $\mathbf{h}_{t}$ serves as a conditional vector and a haze component. This enables the model to learn how clear images and haze gradually combine, allowing it to remove haze during inference physically. Theoretical analyses, including forward process convergence, variational lower bound derivation, and the ELBO formulation, are provided in the appendix for completeness.
3.4 The IDDM Sampling Process
The determinisic implicit sampling process of IDDM can be derived by referring DDIM [DDIM], shown in Fig. 2.
$\mathbf{x_{t-1}}=$ $\sqrt{\bar{\alpha}_{t-1}}\left(\mathbf{x}_{0}+\mathbf{h}_{t}\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\boldsymbol{\epsilon}_{\theta}^{(t)}+\sigma_{t}\boldsymbol{\epsilon}_{\theta}^{(t)}$ (11)
@@ 3.4 The IDDM Sampling Process
where $\boldsymbol{\epsilon}_{\theta}^{(t)}=\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},\mathbf{h}_{t},\mathbf{\hat{x}},t\right)$. Then based on Eq. (11), computational cost can be reduced by utilizing only a subset of timestep indices, which is the same as DDIM [DDIM]. Instead of considering all timesteps from $1$ to $T$, a subsequence $(\tau_{1},\tau_{2},\ldots,\tau_{s})$ is selected. This subsequence is determined by evenly spacing out indices from the complete sequence $(1,2,\ldots,T)$ using the formula $T_{i}=(i-1)\cdot\frac{T}{S}+1$, where $i$ ranges from $1$ to $\tau_{s}$ and $s$ is a parameter controlling the sampling rate. Therefore, for step $\tau_{1}<\tau_{2}$, the sampling process is given by
@@
The predicted $\mathbf{x}_{0}$ from the diffusion process corresponds to the attenuated clear image ($JT_{r}$), while the haze component $\mathbf{h}_{T}$ is utilized to infer the transmission map $T_{r}$. The final clear image $J$ is recovered based on the Atmospheric Scattering Model (ASM) as $J=\mathbf{x}_{0}/T_{r}$. Crucially, to simplify the restoration and avoid the need for explicit atmospheric light ($A^{\prime}$) estimation, we leverage the property that the haze component $\mathbf{h}_{T}$ is proportional to $A^{\prime}$. We therefore implicitly derive the transmission map $T_{r}$ by normalizing the estimated $\mathbf{h}_{T}$ at the final stage. The final restoration is computed as:
$J=\frac{\mathbf{x}_{0}}{1-\mathbf{\tilde{h}}_{T}},$ (14)
where $\mathbf{\tilde{h}}_{T}$ is the stabilized and normalized haze component, representing $\mathbf{h}_{T}/A^{\prime}$. This $\mathbf{\tilde{h}}_{T}$ is obtained by first applying a Gaussian filter to smooth the predicted $\mathbf{h}_{T}$, and then normalizing its pixel values to $[0,1]$. This procedure effectively decouples the haze component from $A^{\prime}$ and leads to superior visual quality. Prior multi-component estimation approaches [Varghese2023] often require dedicated networks or separate modules to estimate $T_{r}$, $A$, and their combination individually. Conversely, we simplifies the restoration task by focusing solely on the direct synthesis of the haze component $\mathbf{h}_{t}$, contributing to a more efficient and robust restoration process.
@@ 4 Method
Figure 3: Visual comparisons on dense hazy images from RTTS dataset [reside] and Fattal’s dataset [fattal]. All models are trained on OTS dataset. Please zoom in for more details.
Figure 4: Visual comparisons on a low-light hazy image from RTTS dataset [reside]. All models are trained on ITS dataset.
Our training framework implements a two-stage sequential strategy that leverages the physical principles of IDDM’s forward process. This design enables the network to jointly perform physics-guided dehazing and distribution-aware denoising during inference.
4.1 Stage I: Physics-guided Denoising Training
The first stage trains a denoising U-Net $\boldsymbol{\tilde{\epsilon}}_{\theta}$ to predict noise $\boldsymbol{\epsilon}_{t}$ from the diffused state $\mathbf{x}_{t}$, conditioned on both the ground-truth haze component $\mathbf{h}_{t}$ and the target hazy image $\hat{\mathbf{x}}$ (Fig. 2a). The training objective follows the standard diffusion loss:
$\mathcal{L}_{s_{1}}=\mathbb{E}_{t\sim[1,T],\mathbf{x}_{0}\sim q\left(\mathbf{x}_{0}\right)}\left\|\boldsymbol{\epsilon}_{t}-\boldsymbol{\tilde{\epsilon}_{\theta}}\left(\mathbf{x}_{t},\mathbf{h}_{t},\hat{\mathbf{x}},t\right)\right\|^{2},$ (15)
where $\mathbf{x}_{t}$ is computed via the IDDM forward process (Eq. (8)). By conditioning on $\mathbf{h}_{t}$, the model learns to denoise under varying physical haze densities, establishing the foundation for physics-aware restoration.
4.2 Stage II: Diffusion-based Haze Estimator Training
For real-world degraded images where ground-truth $\mathbf{h}_{t}$ is unavailable, we train a lightweight haze estimator (HtNet) to predict $\tilde{\mathbf{h}}_{t}$ from the input $\tilde{\mathbf{x}}$ at any timestep $t$. HtNet shares the U-Net architecture of Stage I but with reduced capacity (1/3 input channels).
Critically, HtNet training is supervised by both ground-truth haze and the frozen Stage I network $\boldsymbol{\tilde{\epsilon}}_{\theta}$, creating a bi-directional enhancement mechanism (Fig. 2c). The dual-constraint objective ensures: (1) predicted haze matches ground-truth, and (2) noise prediction using estimated haze aligns with that using ground-truth haze:
This formulation enables HtNet to learn physically meaningful haze representations across all timesteps, while the frozen diffusion model acts as a consistency regularizer. By learning timestep-dependent haze concentrations, the model progressively removes haze over $T$ iterations during inference, naturally aligning with the physical accumulation process.
Training Dataset OTS ITS
@@ 4.2 Stage II: Diffusion-based Haze Estimator Training
Dehamer [dehamer] 33.8582 4.9134 46.5483 4.6256 29.3234 4.5819 45.7394 4.5998
DEA-Net [chen2023dea] 31.1197 4.9206 48.1424 4.6499 29.7903 4.8192 45.813 4.6635
MB-TaylorFormer [mbtaylor] 32.6807 4.8664 46.8968 4.6184 32.8450 4.9090 47.7943 4.6302
C2PNet [c2pnet] 34.2951 5.0336 48.6310 4.6399 33.7423 5.0546 48.5107 4.6739
PSD [psd] 25.2394 3.8742 30.5665 4.3459 N/A N/A N/A N/A
D4 [D4] N/A N/A N/A N/A 33.206 5.8548 57.7313 3.7239
OKNet [OKNET] 35.1734 5.0858 49.1408 4.6548 33.1591 4.9664 49.1645 4.6558
IDDM 18.3210 3.4937 27.0719 4.8401 27.8151 4.2198 38.9139 4.9153
Table 1: The quantitative results of our method and the state-of-the-arts on real-world haze dataset RTTS. The best results are shown in bold.
Method OTS ITS
[…]
Dehamer 0.795 0.950 0.909 0.805 0.942 0.909
DEA-Net 0.882 0.973 0.949 0.840 0.934 0.899
MB-TaylorFormer 0.793 0.933 0.898 0.793 0.933 0.898
C2PNet 0.880 0.970 0.948 0.795 0.913 0.874
D4/PSD 0.860 0.966 0.925 0.814 0.951 0.923
OKNet 0.875 0.973 0.947 0.846 0.935 0.908
IDDM (Ours) 0.903 0.973 0.956 0.879 0.959 0.936
Table 2: Quantitative comparison on BEDDE dataset. The best results are shown in bold.
Method BEDDE RTTS
VI RI VSI Brisque NIQE PIQE NIMA
RDDM 0.778 0.957 0.908 22.391 6.876 43.704 4.270
Diff-Plugin 0.785 0.956 0.908 25.651 4.613 28.339 4.915
RIDCP 0.864 0.967 0.938 18.782 4.157 28.254 4.427
IPC-Dehaze 0.877 0.970 0.943 19.447 4.084 31.644 4.865
Ours 0.878 0.979 0.946 17.734 3.7637 26.771 4.929
Table 3: Comparison with two diffusion-based methods and two real-world dehazing methods on BEDDE and RTTS datasets. All methods are trained on the RIDCP dataset.
@@ 5.1 Experimental Settings
Training Datasets. During training, we use ITS (13,990 indoor synthetic hazy image pairs) and OTS (72,135 outdoor synthetic hazy image pairs) datasets from RESIDE [reside]. ITS depth maps come from NYU Depth Dataset V2 [Silberman:ECCV12], while OTS provides its own depth maps. We implement an online data generation strategy during training. For each clear image and corresponding depth map, we synthesize both the hazy image $\hat{\mathbf{x}}$ and the corresponding $\mathbf{h}_{t}$ at time step $t$ using the Atmospheric Scattering Model (ASM). Parameters such as airlight $A\in[0.7,1.0]$ and scattering coefficient $\sigma\in[0.4,1.5]$ are randomly sampled to guarantee data diversity and robustness across various haze conditions.
Testing Datasets. We qualitatively and quantitatively evaluate our method against state-of-the-art techniques on RTTS [reside] (4,322 real-world hazy images) and BEDDE [Bedde] (209 paired real-world images) datasets, with additional visual comparisons using Fattal’s dataset [fattal].
Implementation Details. During the training stages, experiments are conducted on 4 NVIDIA GeForce RTX 4090 GPUs. We optimize our networks with Adam optimizer ($\beta_{1}=0.9,\beta_{2}=0.99$). The learning rate is set to 0.00001 and the batch size is set to 16. For data augmentation, we apply random horizontal and vertical flips, crop images to 256×256 pixels, and randomly introduce JPEG compression artifacts, gamma adjustment, and colorful haze based on [Wu_2023_CVPR]. At stages 1, the diffusion denoising Unet is trained for 500,000 iterations, with the parameter $T$ is set as 1000. At stage 2, HtNet is trained for 60,000 iterations. Finally, all the evaluations for all methods are conducted on a NVIDIA GeForce RTX 3090 GPU.
(a) Input
(b) Diff-Plugin
(c) IPC-Dehaze
(d) RIDCP
(e) RDDM
(f) Ours
Figure 5: Visual comparison of methods from Table. 3 on a very dense hazy image from RTTS dataset [reside].
@@ 5.2 Results and Comparisons
Quantitative Comparison. To evaluate the robustness of domain generalization, we train our models on both indoor and outdoor synthetic hazy datasets from RESIDE [reside] and test them on two real-world datasets: RTTS [reside] and BEDDE [Bedde]. We conduct comprehensive comparisons with state-of-the-art approaches: end-to-end deep learning methods, (Dehamer [dehamer], DEA-Net [chen2023dea], MB-TaylorFormer [mbtaylor], C2PNet [c2pnet], PSD [psd], D4 [D4], OKNET [OKNET]), two real-world image dehazing methods which employ different training datasets beyond the standard ITS or OTS (RIDCP [Wu_2023_CVPR], IPC-Dehaze [IPCDehaze]), and two diffusion-based approaches (Diff-Plugin [diffplugin], RDDM [RDDM]). Notably, driven by the requirements of their methods, PSD [psd] is trained exclusively on outdoor hazy images, while D4 [D4] is trained solely on indoor hazy images. As shown in Tables 1, 2, and 3, the experimental results demonstrate that our IDDM consistently outperforms all competing approaches, achieving the best performance across seven evaluation metrics.
Qualitative Comparison. As shown in Fig. 3 and Fig. 4, we compare the visual results of our model with previous state-of-the-art methods on the RTTS dataset [reside] and Fattal’s data [fattal]. In Fig. 3, all models are trained on outdoor synthetic dataset, while in Fig. 4, all models are trained on the indoor synthetic dataset. As shown in Fig. 4, even when the training and test sets span two domains (indoor to outdoor, synthetic to real), IDDM still shows bright and clear dehazing performance. In contrast, the other methods tend to produce dark tones in the dehazed images, likely due to being trained exclusively on indoor hazy images.
@@
(d) w/o Eq. 14
(e) ours (OTS)
(f) ours (ITS)
Figure 6: Visual ablation results of the IDDM.
@@
$\sigma\in[0.1,1.0]$ 0.641 0.948 0.838 $T=300$ 0.669 0.905 0.865
$\sigma\in[0.4,1.5]$ 0.879 0.959 0.936 $T=500$ 0.750 0.922 0.881
$\sigma\in[1.0,2.0]$ 0.616 0.911 0.822 $T=1000$ 0.879 0.959 0.936
Table 5: Ablation studies on hyperparameters $\sigma$ (scattering coefficient range) and $T$ (diffusion timesteps) evaluated on the BEDDE dataset. Models are trained on OTS.
@@
Table 6: Ablation studies on hyperparameter AA (atmospheric light range) evaluated on the BEDDE dataset. Models are trained on ITS.